# NegotiationToM: A Benchmark for Stress-testing Machine Theory of Mind on Negotiation Surrounding

### v1

@@ Abstract
Large Language Models (LLMs) have sparked substantial interest and debate concerning their potential emergence of Theory of Mind (ToM) ability. Theory of mind evaluations currently focuses on testing models using machine-generated data or game settings prone to shortcuts and spurious correlations, which lacks evaluation of machine ToM ability in real-world human interaction scenarios. This poses a pressing demand to develop new real-world scenario benchmarks. We introduce NegotiationToM, a new benchmark designed to stress-test machine ToM in real-world negotiation surrounding covered multi-dimensional mental states (i.e., desires, beliefs, and intentions). Our benchmark builds upon the Belief-Desire-Intention (BDI) agent modeling theory and conducts the necessary empirical experiments to evaluate large language models. Our findings demonstrate that NegotiationToM is challenging for state-of-the-art LLMs, as they consistently perform significantly worse than humans, even when employing the chain-of-thought (CoT) method.
@@ 1 Introduction
Theory of Mind (ToM) was introduced as an agent’s capacity to infer the mental states of others, such as desires, beliefs, and intentions Premack and Woodruff (1978); Ma et al. (2023). Numerous scenarios involving human cognition and social reasoning rely on the ToM modeling of others’ mental states Gopnik and Wellman (1992); Baron-Cohen (1997); Gunning (2018), such as comprehending and forecasting others’ actions Dennett (1988), planning over others’ beliefs and subsequent actions Favier et al. (2023), and various forms of reasoning and decision-making Pereira et al. (2016); Rusch et al. (2020). However, the traditional evaluation benchmarks for language models are primarily theoretical game settings or synthetic template-based data generated by the large language model, which may inherently suffer from shortcuts and spurious correlations Sclar et al. (2023); Ullman (2023); Shapira et al. (2023a); Ma et al. (2023). Consequently, these benchmarks assess language models from a theoretical perspective, which may not precisely and effectively reflect the ToM capabilities of large language models in practical situations.
In reality, ToM ability plays a crucial role in comprehending dynamic social interactions (e.g., negotiation conversations) by forming an essential element of effective communication Frith (1994); Schober (2005), and inferring other’s mental states in a conversation requires machines as humans to comprehend text beyond surface forms of utterance and utilize the incomplete information presented in the conversation. ToM is closely related to interpersonal social intelligence Ganaie and Mudasir (2015); Stone (2006); Williams et al. (2022); Sap et al. (2022), which allows us to navigate and understand social situations ranging from simple everyday interactions to complex negotiations Yang et al. (2021); de Weerd et al. (2017); Gardner et al. (1995); de Weerd et al. (2013).
@@ 3.1 Design Considerations for NegotiationToM
There are several essential design considerations we go through when constructing the NegotionToM. (1) the scenario of the dataset should be grounded in a human-to-human real-world negotiation (e.g., real-world camping scenario). (2) the dataset should be a natural conversational dataset instead of generated from a synthetic template to avoid reporting bias Gordon and Durme (2013) and shortcuts Aru et al. (2023). (3) the dataset should be equipped with abundant and diverse linguistic features and semantic context (e.g., negotiation argument) instead of bargaining on the numerical value or meaningless counter-offer Lewis et al. (2017); He et al. (2018).
@@ 3.2 CaSiNo
Inspired by several considerations above, CaSiNo Chawla et al. (2021b) was employed as the source data and modified to construct NegotiatonToM. CaSiNo is a bilateral human-to-human natural conversational dataset that covers rich linguistic features and many realistic aspects of negotiations, such as small talk, preference elicitation, emotional expression, and convincing strategies based on individual desire. In this dataset, the participants take the role of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. For each conversation, participants discuss individual needs by making various convincing arguments from their camping experiences, such as Personal Care, Recreational, Group Needs, or Emergency requirements. One example of Group Needs is "I need more firewood due to having several people join on the trip and needing a bigger fire overall." We illustrate some of these arguments in Table 21 in Appendix A.2. Therefore, crafting our benchmark from the CaSiNo offers a range of scenarios based on how to align the preferences of the two parties to reveal more interesting behaviors.
@@ 3.3 Theory of Mind in NegotiationToM
Belief refers to a mental state in which an individual assumes a specific stance, attitude, or opinion toward a proposition. In contemporary discussions within the field of philosophy of mind, the term “belief” is commonly employed to denote an individual’s cognitive stance or acceptance of something as true or holding it to be the case Turiel (2008). Note that this notion of belief does not inherently require active reflection, nor does it necessitate truthfulness Armstrong (1973); Moses (1993). In NegationiToM, understanding the state of the opponent’s item preference order, which is explicitly or implicitly expressed in the conversation, is the main way to form the belief. Therefore, the belief question will query the LLMs on what one participant thinks of another participant’s item preferences, given the current round of dialogue with previous conversation history.
@@ 3.4 Annotation & Statistics
NegotiationToM is annotated based on a multi-turn negotiation dialogue corpus, the CaSiNo Chawla et al. (2021a) dataset. Each instance in CaSiNo is an -round alternating dialogue between two participants, and 111When the dialogue ends with user , is an empty utterance.. They take on the roles of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. We adopt the subset with strategy annotations and undertake the annotation on the desire and belief states behind each utterance..
@@
To ensure the annotation quality, we evaluate the workers on the first 100 rounds of conversations and explain to them in detail about their typical errors. More details of the annotation process are in Appendix A.1. Although annotating NegotiationToM requires understanding complex dialogues in CaSiNo, we observed high inter-annotator agreement (Fleiss’s =79.03% Fleiss (1971) for NegotiationToM benchmark, the breakdown computation of are shown in Table 3).
@@ 4.1 Baseline Models
In this work, we test six recent instruction-tuned large language models: GPT-4 OpenAI (2023), ChatGPT (OpenAI, 2022), Claude-v1.3222https://www.anthropic.com/news/introducing-claude Anthropic (2023a), Claude-v2.1333https://www.anthropic.com/news/claude-2-1 Anthropic (2023b), Llama-2 Chat 13B Touvron et al. (2023), and Llama-2 Chat 70B Touvron et al. (2023). Descriptions for each model are in Appendix B.1. By following the common practices in the theory of mind field Kim et al. (2023); Gandhi et al. (2023); Shapira et al. (2023b), we test these models with two types of prompts: (1) one is zero-shot prompting and we utilize the prompt template in Robinson and Wingate (2023) to formulate the task as a multiple choice question answering problem as a baseline. (2) another one is the Chain-of-Thought (CoT) prompting method by following Kojima et al. (2022) and using the prompt “let’s think step by step.” Apart from these two settings, we also assess the LLMs’ performance by using the few-shot setting to validate whether LLMs can improve their performance with input-output exemplars. We append four tailored exemplars for desire and belief states and seven designed exemplars for intentions, covering all the input-output exemplars’ labels. More configuration details can be found in Appendix B.2, and the prompt template refers to Appendix B.4. To measure the specific performance gap between humans and the state-of-the-art machine on the NegotiationToM, we employ three graduate students in computer science to complete the human evaluation task. More details of human evaluation are shown in Appendix C.2.
@@ 5.1 Main Result
Table 4 summarizes the main results of the state-of-the-art large language models in NegotiationToM, from which we derive the following conclusions. First, the performance of all models is significantly worse than human performance, even after employing the zero-shot chain-of-thought (CoT) method. There is a significant performance gap between machines and humans in the ToM negotiation evaluation scenario. Specifically, compared with the best state-of-the-art models in each mental state, the performance gap is 27.85% in desire, 32.96% in belief, 43.82% in Micro F1 score, and 48.98% in Macro F1 score in intention. Second, GPT-4 0613 (CoT) achieves the best performance among all the models regarding inferring desire and belief, while Claude-v2.1 (CoT) outperforms all other models in the intention classification task in NegitiationToM. Third, we observe that most models received an improvement in scores when the chain-of-thought (CoT) method was applied. Nevertheless, there are still significant score gaps compared to human performance.
@@ B.2 Hyperparameter
We use default hyperparameters for all the large language models mentioned in this paper. For ChatGPT (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-0613), the default parameters444https://platform.openai.com/docs/api-reference/chat/create are temperature=1 and top_p=1. For LLaMa2-Chat(13B) and LLaMa2-Chat(70B) models, we follow the default setting where temperature=0.5, top_p=0.9. For Claude-v1.3 and Claude-v2.1, the default parameters are temperature=0.5, top_p=1.

### v2

@@ Abstract
Large Language Models (LLMs) have sparked substantial interest and debate concerning their potential emergence of Theory of Mind (ToM) ability. Theory of mind evaluations currently focuses on testing models using machine-generated data or game settings prone to shortcuts and spurious correlations, which lacks evaluation of machine ToM ability in real-world human interaction scenarios. This poses a pressing demand to develop new real-world scenario benchmarks. We introduce NegotiationToM111The dataset is available at https://github.com/HKUST-KnowComp/NegotiationToM, a new benchmark designed to stress-test machine ToM in real-world negotiation surrounding covered multi-dimensional mental states (i.e., desires, beliefs, and intentions). Our benchmark builds upon the Belief-Desire-Intention (BDI) agent modeling theory and conducts the necessary empirical experiments to evaluate large language models. Our findings demonstrate that NegotiationToM is challenging for state-of-the-art LLMs, as they consistently perform significantly worse than humans, even when employing the chain-of-thought (CoT) method.
@@ 1 Introduction
Theory of Mind (ToM) was introduced as an agent’s capacity to infer the mental states of others, such as desires, beliefs, and intentions Premack and Woodruff (1978); Ma et al. (2023). Numerous scenarios involving human cognition and social reasoning rely on the ToM modeling of others’ mental states Gopnik and Wellman (1992); Baron-Cohen (1997); Gunning (2018), such as comprehending and forecasting others’ actions Dennett (1988), planning over others’ beliefs and subsequent actions Favier et al. (2023), and various forms of reasoning and decision-making Pereira et al. (2016); Rusch et al. (2020). Some previous research believes that LLMs already exhibit a high level of competence in addressing ToM tasks Strachan et al. (2024); Bubeck et al. (2023); Kosinski (2023), while other studies express doubt and develop benchmarks to illustrate that LLMs do not possess proficient ability in ToM tasks Sap et al. (2022); Ullman (2023); Shapira et al. (2024). However, these traditional evaluation benchmarks for language models are primarily theoretical game settings or synthetic template-based data generated by the large language model, which may inherently suffer from shortcuts and spurious correlations Sclar et al. (2023); Ullman (2023); Shapira et al. (2023a); Ma et al. (2023). Consequently, these benchmarks assess language models from a theoretical perspective, which may not precisely and effectively reflect the ToM capabilities of large language models in practical situations.
In reality, ToM ability plays a crucial role in comprehending dynamic social interactions (e.g., negotiation conversations) by forming an essential element of effective communication Frith (1994); Schober (2005), and inferring other’s mental states in a conversation requires machines as humans to comprehend text beyond surface forms of utterance and utilize the incomplete information presented in the conversation. ToM is closely related to interpersonal social intelligence Ganaie and Mudasir (2015); Stone (2006); Williams et al. (2022); Sap et al. (2022), which allows us to navigate and understand social situations ranging from simple everyday interactions to complex negotiations Yang et al. (2021); Kim et al. (2023); de Weerd et al. (2017); Gardner et al. (1995); de Weerd et al. (2013).
@@ 3.1 Design Considerations for NegotiationToM
There are several essential design considerations we go through when constructing the NegotiationToM. (1) the scenario of the dataset should be grounded in a human-to-human real-world negotiation (e.g., real-world camping scenario). (2) the dataset should be a natural conversational dataset instead of generated from a synthetic template to avoid reporting bias Gordon and Durme (2013) and shortcuts Aru et al. (2023). (3) the dataset should be equipped with abundant and diverse linguistic features and semantic context (e.g., negotiation argument) instead of bargaining on the numerical value or meaningless counter-offer Lewis et al. (2017); He et al. (2018).
@@ 3.2 CaSiNo
Inspired by several considerations above, CaSiNo Chawla et al. (2021b) was employed as the source data and modified to construct NegotiationToM. CaSiNo is a bilateral human-to-human natural conversational dataset that covers rich linguistic features and many realistic aspects of negotiations, such as small talk, preference elicitation, emotional expression, and convincing strategies based on individual desire. In this dataset, the participants take the role of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. For each conversation, participants discuss individual needs by making various convincing arguments from their camping experiences, such as Personal Care, Recreational, Group Needs, or Emergency requirements. One example of Group Needs is "I need more firewood due to having several people join on the trip and needing a bigger fire overall." We illustrate some of these arguments in Table 21 in Appendix A.2. Therefore, crafting our benchmark from the CaSiNo offers a range of scenarios based on how to align the preferences of the two parties to reveal more interesting behaviors.
@@ 3.3 Theory of Mind in NegotiationToM
Belief refers to a mental state in which an individual assumes a specific stance, attitude, or opinion toward a proposition. In contemporary discussions within the field of philosophy of mind, the term “belief” is commonly employed to denote an individual’s cognitive stance or acceptance of something as true or holding it to be the case Turiel (2008). Note that this notion of belief does not inherently require active reflection, nor does it necessitate truthfulness Armstrong (1973); Moses (1993). In NegotiationToM, understanding the state of the opponent’s item preference order, which is explicitly or implicitly expressed in the conversation, is the main way to form the belief. Therefore, the belief question will query the LLMs on what one participant thinks of another participant’s item preferences, given the current round of dialogue with previous conversation history.
@@ 3.4 Annotation & Statistics
NegotiationToM is annotated based on a multi-turn negotiation dialogue corpus, the CaSiNo Chawla et al. (2021a) dataset. Each instance in CaSiNo is an -round alternating dialogue between two participants, and 222When the dialogue ends with user , is an empty utterance.. They take on the roles of campsite neighbors and negotiate for food, water, and firewood packages for their upcoming trip. We adopt the subset with strategy annotations and undertake the annotation on the desire and belief states behind each utterance..
@@
To ensure the annotation quality, we evaluate the workers on the first 100 rounds of conversations and explain to them in detail about their typical errors. More details of the annotation process are in Appendix A.1. Although annotating NegotiationToM requires understanding complex dialogues in CaSiNo, we observed high inter-annotator agreement. The Fleiss’s is 79.03% Fleiss (1971) for NegotiationToM benchmark, the breakdown computation of are shown in Table 3.
@@ 4.1 Baseline Models
In this work, we test six recent instruction-tuned large language models: GPT-4 OpenAI (2023), ChatGPT (OpenAI, 2022), Claude-v1.3333https://www.anthropic.com/news/introducing-claude Anthropic (2023a), Claude-v2.1444https://www.anthropic.com/news/claude-2-1 Anthropic (2023b), Llama-2 Chat 13B Touvron et al. (2023), and Llama-2 Chat 70B Touvron et al. (2023). Descriptions for each model are in Appendix B.1. By following the common practices in the theory of mind field Kim et al. (2023); Gandhi et al. (2023); Shapira et al. (2023b), we test these models with two types of prompts: (1) one is zero-shot prompting and we utilize the prompt template in Robinson and Wingate (2023) to formulate the task as a multiple choice question answering problem as a baseline. (2) another one is the Chain-of-Thought (CoT) prompting method by following Kojima et al. (2022) and using the prompt “let’s think step by step.” Apart from these two settings, we also assess the LLMs’ performance by using the few-shot setting to validate whether LLMs can improve their performance with input-output exemplars. We append four tailored exemplars for desire and belief states and seven designed exemplars for intentions, covering all the input-output exemplars’ labels. More configuration details can be found in Appendix B.2, and the prompt template refers to Appendix B.4. To measure the specific performance gap between humans and the state-of-the-art machine on the NegotiationToM, we employ three graduate students in computer science to complete the human evaluation task. More details of human evaluation are shown in Appendix C.2.
@@ 5.1 Main Result
Table 4 summarizes the main results of the state-of-the-art large language models in NegotiationToM, from which we derive the following conclusions. First, the performance of all models is significantly worse than human performance, even after employing the zero-shot chain-of-thought (CoT) method. There is a significant performance gap between machines and humans in the ToM negotiation evaluation scenario. Specifically, compared with the best state-of-the-art models in each mental state, the performance gap is 27.85% in desire, 32.96% in belief, 43.82% in Micro F1 score, and 48.98% in Macro F1 score in intention. Second, GPT-4 0613 (CoT) achieves the best performance among all the models regarding inferring desire and belief, while Claude-v2.1 (CoT) outperforms all other models in the intention classification task in NegotiationToM. Third, we observe that most models received an improvement in scores when the chain-of-thought (CoT) method was applied. Nevertheless, there are still significant score gaps compared to human performance.
@@ B.2 Hyperparameter
We use default hyperparameters for all the large language models mentioned in this paper. For ChatGPT (gpt-3.5-turbo-0613) and GPT-4 (gpt-4-0613), the default parameters555https://platform.openai.com/docs/api-reference/chat/create are temperature=1 and top_p=1. For LLaMa2-Chat(13B) and LLaMa2-Chat(70B) models, we follow the default setting where temperature=0.5, top_p=0.9. For Claude-v1.3 and Claude-v2.1, the default parameters are temperature=0.5, top_p=1.